If a sprinter starts from rest and reaches a speed of 3.3 m/s after 1.5s, how far did they travel is those 1.5s?
1
Expert's answer
2018-02-16T09:52:08-0500
Acceleration of the sprinter is 3.3/1.5 = 2.2 m/s2
Again we know,
S = 1/2 f〖 t〗^2 ……………………………..(1)
Where , initial velocity = 0, S = distance, f= acceleration, t= time
Here, f = 2.2 m/s2, t = 1.5 second
Now put these values in equation 1, we get,
S =1/2 x 2.2 x 〖 1.5〗^2 = 2.475 metre.